OpenStudy (anonymous):

20 + 25 + 50(x - 2) = C Arianna paid a total of $145 to check all her bags on a flight. How many bags did she check in all?

OpenStudy (anonymous):

20 + 25 + 50(x - 2) = C 20 + 25 + 50(x - 2) = 145 can you solve it

OpenStudy (lukebluefive):

$145 - ($20 + $25) = $50(x - 2) $145 - ($45) = $50(x - 2) $100 = $50(x - 2) Just divide each side by $50, add two to both sides, and you have your answer.